OpenStudy (anonymous):

The length of a rectangle is 5 in more than its width. The perimeter of the rectangle is 42 inches. Find the WIDTH of the rectangle.

OpenStudy (anonymous):

let width be x

OpenStudy (anonymous):

then length will be x+5

OpenStudy (anonymous):

perimeter is 2(l+b) so 2(x+x+5)=42

OpenStudy (anonymous):

2(2x+5)=42

OpenStudy (anonymous):

x=21-5/2

OpenStudy (anonymous):

16/2=8

OpenStudy (anonymous):

so length is 8 +5=13

OpenStudy (anonymous):

thanks